OpenStudy (magbak):

Factor completely p^4 – 8p^2 + 16 (p^2 – 4)^2 (p – 2)(p + 2)(p^2 + 4) (p – 2)^2(p + 2)^2 (p – 2)(p + 2)(p^2 – 4)

OpenStudy (nurali):

p^4 - 8p^2 + 16 (p^2 - 4)^2 [(p - 2)(p + 2)]^2 (p - 2)^2 (p + 2)^2

OpenStudy (magbak):

So it is c

OpenStudy (nurali):

Yup

OpenStudy (jhannybean):

\[\large a^2-b^2 = (a+b)(a-b)=(a-b)^2=a^2-2ab+b^2\]\[\large a^2 = p^4 \ \therefore \ a=\sqrt{p^4}=p^2\]\[\large b^2 = 16 \ \therefore \ b=\sqrt{16}=4\]Putting it into form \[\large (p^2-4)^2\]\[\large a^2 = p^2 \ \therefore \ a=p\]\[\large b^2= 4 \ \therefore \ b=2 \]\[\large [(p+2)(p-2)]^2\]distribute the square. \[\large (p+2)^2(p-2)^2\]